Question 4 Bakesale Enterprises purchased equipment on May 1, 2018 for $6,300. The company expects to use the equipment for 5What adjusting journal entry should the company make at the end of each month if monthly financials are prepared (annual depreciation is $1,260)? (Credit account titles are automatically indented when the amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)
Answer #1

Ans:

S.no Account title and explanation Debit($) Credit($)
1. Depreciation expense 105
Accumulated depreciation 105
( to record adjusting entry )

Yearly monthly depreciation= 6,300/5

= 1,260

Monthly depreciation= 1,260/12

=105

2. Book value of depreciation= Cost - Accumulated depreciation

= 6,300-105

= 6,195